HUBBLE SPACE TELESCOPE DAILY REPORT #2145 PERIOD COVERED: 0000Z (UTC) 06/15/98 - 0000Z (UTC) 06/16/98 Daily Status Report as of 167/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Seven Sets of STIS/CCD 8057 (STIS CCD Performance Monitor {continued}) The Space Telescope Imaging Spectrograph (CCD) was used to monitor to overall CCD performance. During the final iteration of this proposal, the two losses of lock mentioned in 2.1 and HSTAR 6541 occurred. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the Kuiper Belt object Pholus in order to study its surface composition. The surfaces of Kuiper Belt Objects contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Four Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies IC5052, NGC-1291, NGC-3034 and NGC-2441 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.5 Completed Two Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. During the final iteration of this proposal, the two losses of lock mentioned in 2.1 and HSTAR 6541 occurred. In this iteration also was another loss of lock, mentioned in HSTAR 6542, at 166/121120Z. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. During the final iteration of this proposal, the two losses of lock mentioned in 2.1 and HSTAR 6541 occurred. In this iteration also was another loss of lock, mentioned in HSTAR 6542, at 166/121120Z. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.7 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed NIC/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of M-Class dwarf star GJ748AB. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/1/2/3 7822 (Grism Spectra of Centaurs and TNOs from 1-2 Microns) The Near Infrared Camera (Cameras One, Two and Three) was used to make grism spectra of object 1995DW2 from 1-2 microns. The surfaces of Centaurs and Trans-Neptunian Objects (TNOs) contain potentially unique records of early evolutionary processing {accretional and dynamical} in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.11 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.12 Completed NIC/3 7821 (SINS: The Supernova Intensive Study - Cycle 7+: Scarlet SINS) The Near Infrared Camera (Camera Three) was used to make observations of the super nova LMC-SN1987 which is located in the Large Magellanic Cloud. Supernovae create the chemical history of the Universe, energize the interstellar gas, and stand at the center of efforts to understand the extragalactic distance scale. During this proposal, losses of lock (defined in HSTAR 6542) occurred at 166/210145Z, 166/223828Z, 166224003Z, 166/224117Z, 167/001615Z, 167/002856Z, 167/015154Z, 167/020538Z, 167/021819Z, 167/022837Z, 167/033011Z, 167/034222Z, and 167/035504Z. The observations were completed with no further problems noted. 1.13 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. During this proposal, losses of lock (defined in HSTAR 6542) occurred at 166/210145Z, 166/223828Z, 166224003Z, 166/224117Z. Otherwise, the observations were completed as scheduled, and no further problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 9 Successful : 9 HSTAR 6541 was written to document two losses of lock following a type 3 slew at 166/1032Z. The vehicle recovered fine lock both times. The proposals described in 1.1, 1.5 and 1.6 were affected by these losses of lock. HSTAR 6542 was written to document multiple losses of lock for the period 166/1000Z - 167/1000Z. Affected proposals in this operational day are described in 1.5, 1.6, 1.12, and 1.13 above, along with the number and time of the pertinent losses of lock. Other affected proposals, which occur during the period of tomorrow's report, will be reported then. 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: The NSSC-1 status buffer was dumped at 166/0943Z using ROP NS-3. At 166/114013Z, the STIS MCE-1 reset outside an SAA event and while the MAMA-1 low voltage was on. The MAMA-1 will be recovered via the SMS at 168/034138Z, the time of the next MAMA-1 high voltage on. HSTAR 6540 was written. STIS science operations were not affected. The SSR EDAC error counter was cleared two times using ROP SR-1. As directed by ROP DF-7, the DF244 ephemeris tables were uplinked ar 167/0030Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH